In a batch of 15 jars of instant coffee, 5 have been under filled. Suppose 6 of these jars are selected without replacement. If more than 1 of these selected jars are under filled, the whole batch is sent back for refill. What is the probability of the batch being sent back for refill? (Round off to 2 decimal places.)

1
Expert's answer
2022-05-04T10:52:31-0400

Suppose that "X" is a random variable and it denotes a number of selected under filled jars. The aim is to find the probability "P(X>1)". "P(X>1)=1-P(X=0)-P(X=1)". There are "C_{15}^6=\frac{15!}{9!6!}=\frac{10\cdot11\cdot12\cdot13\cdot14\cdot15}{6!}=5005" different ways to choose "6" jars from "15". There are "C_{10}^6=\frac{10!}{4!6!}=\frac{7\cdot8\cdot9\cdot10}{4!}=210" ways to choose "6" filled jars from "10". There are "C_{10}^5=\frac{10!}{5!5!}=\frac{6\cdot7\cdot8\cdot9\cdot10}{5!}=252" ways to get "5" filled jars from "10". There are "5" ways to select "1" under filled jar from "5". "P(X=0)=\frac{C_{10}^6}{C_{15}^6}=\frac{210}{5005}\approx0.042". "P(X=1)=\frac{5\cdot C_{10}^5}{C_{15}^6}=\frac{5\cdot210}{5005}\approx0.210". Thus, the probability that the batch will be sent back is: "1-0.042-0.210\approx0.748".

Answer: the probability that the batch will be sent back is: "0.75" (it is rounded to "2" decimal places).